OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (anonymous):

f(x)= -4(10x-6) +7 =-40x+24+7 =-40x+31

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

no problem ^.^